OBSERVATION OF THE Ã^{2}A′ - X̃^{2}A′′ ELECTRONIC TRANSITION OF VINOXY RADICAL USING CAV-ITY RINGDOWN SPECTROSCOPY

\underline{ERIN N. SHARP}, GABRIEL M.P. JUST, AND TERRY A. MILLER, Laser Spectroscopy Facility, Department of Chemistry, The Ohio State University, Columbus, OH 43210.

Free radicals are key components in the oxidation of hydrocarbons, both in combustion and in our atmosphere. More specifically, the vinoxy radical, CH_{2}CHO, is a prototypical alkenoxy radical, which is known to play an important role in the reaction of O(^{3}P) and OH with olefins and olefinic radicals. While the B̃ - X̃ transition of vinoxy has been studied in considerable detail over the last twenty years, the Ã - X̃ transition remains relatively unexplored because of its much weaker absorption cross-section. Due to its low oscillator strengh, cavity ringdown spectroscopy (CRDS) has been applied to study the Ã - X̃ near-IR electronic transition of vinoxy radical. In addition, ab initio calculations were conducted in order to predict the Ã - X̃ origin frequency, as well as to aid in assigning other vibrational structure in the spectrum.